# Interactive Identification of Granular Materials using Force Measurements

### v1

@@ Abstract
The ability to identify granular materials facilitates the emergence of various new applications in robotics, ranging from cooking at home to truck loading at mining sites. However, granular material identification remains a challenging and underexplored area. In this work, we present a novel interactive material identification framework that enables robots to identify a wide range of granular materials using only a force-torque sensor for perception. Our framework, comprising interactive exploration, feature extraction, and classification stages, prioritizes simplicity and transparency for seamless integration into various manipulation pipelines. We evaluate the proposed approach through extensive experiments with a real-world dataset comprising 11 granular materials, which we also make publicly available. Additionally, we conducted a comprehensive qualitative analysis of the dataset to offer deeper insights into its nature, aiding future development. Our results show that the proposed method is capable of accurately identifying a wide range of granular materials solely relying on force measurements obtained from direct interaction with the materials. Code and dataset are available at: https://irobotics.aalto.fi/indentify_granular/.
@@ I Introduction
The ability to accurately identify granular materials holds great potential across various domains of robotics, paving the way for future advancements in assistive robotics as well as autonomous working machines. In assistive robotics, being able to distinguish different food ingredients from each other is vital for planning and performing food-related downstream tasks such as assistive feeding or assistive cooking. Similarly, in industries like mining or construction, autonomous working machines must possess the capacity to differentiate between different granular materials for tasks such as truck loading and unloading or soil cultivating. Furthermore, material identification serves beyond mere manipulation tasks; it can also be employed for anomaly detection and quality assurance in production lines, contributing to overall efficiency and reliability.
Prior works on granular material identification rely either on tactile array sensors or audio sensing and identify materials based on their vibration ”signature” when the material is shaken in a closed container [eppe2018deep, jin2019open, chen2016learning, wang2023stev]. However, the lack of direct interaction means that the models’ classification decisions are based on material properties that may not best correspond to the material dynamics during manipulation, making the approaches suboptimal for use as a part of a manipulation pipeline.
To address this gap, we introduce a novel interactive perception framework that identifies granular materials by interacting with them directly. Specifically, our method measures feedback from the interaction with a standard 6-axis force-torque sensor (F/T sensor) and subsequently classifies the materials based on the features extracted from the raw F/T measurements. The F/T sensor is mounted on the robot’s end effector, which is partially submerged in the material during a predefined motion. We opt for a pure F/T sensor approach due to its simplicity and widespread availability of cost-effective sensors, ensuring ease of implementation across diverse real-world settings. The proposed approach enables robots to accurately identify the materials they interact with, enhancing their ability to understand and respond effectively to the environment. To the best of the authors’ knowledge, this work is the first attempt to identify granular materials using only a regular F/T sensor for perception. The contributions of this work include:
  • A novel interactive perception framework powered with a classification model tailored for granular material identification.
  • A novel real-world dataset of force-torque measurements containing 62 samples out of 11 different materials, totaling 682 individual samples.
  • A comprehensive analysis of the dataset, facilitates a deeper understanding of its underlying characteristics, the distinguishability of materials, and identification of their distinctive features.
@@ II-A Interactive perception with 6-axis F/T sensor
The majority of works on interactive perception with F/T sensors focus on identifying surface materials [liu12] or estimating surface properties [le2021probabilistic] of rigid bodies. Typically, these studies involve attaching the F/T sensor to an artificial fingertip or similar end-effector, which interacts with the material surface to explore its properties. For example, Lam et al. [lam2014study] use this approach to identify 18 different surface materials. The authors reduce the raw data from the interaction into feature vectors of 3, 4, or 5 time-domain features and compare different classifiers, such as the k-nearest neighbor method, support vector machines, and neural networks, on their ability to distinguish the materials. Similarly, the work from Markert et al. [markert2021fingertip] proposed an interaction procedure for an artificial fingertip, which first gathers F/T data from 21 different surface materials, then classifies the materials with high accuracy using features derived from the F/T measurements’ spectral distributions. Furthermore, Le et al. [le2021probabilistic] integrated F/T sensing with visual information to estimate friction maps for multi-material rigid objects.
In addition, F/T sensors have also been used for object recognition and property estimation for rigid objects. For example, Higy et al. [higy2016combining] used 6-DoF F/T sensors combined with robot joint positions to perform haptic classification of rigid objects. However, information from multiple sensor modalities must be combined to achieve reasonably high classification accuracy, as poor performance is observed when each sensor operates independently. In [pavlic2023robotscale], the authors proposed a method that employs a 6-axis F/T sensor attached to the robot’s two-fingered end effector to estimate the mass and inertia of rigid objects.
Prior research highlights that F/T sensors can provide valuable information for various robotic perception tasks. Our work expands the scope of F/T sensing in robotics by introducing a novel application area focusing on deformable objects, in specific, granular materials. Additionally, we present an interpretable method for processing F/T sensor data from interaction with granular materials, further enhancing the applicability of F/T sensing in this domain.
@@ II-B Granular Material Identification
The existing literature on granular material identification is sparse, with only a few initiatives focusing on identifying granular materials in closed containers. Some of those initiatives utilized audio recordings [eppe2018deep, jin2019open], while others leveraged vibration measurements [chen2016learning, wang2023stev]. In addition to identification, some works, such as [matl2020inferring] and [guo2023estimating], discuss the parameter estimation of granular materials.
In audio-based identification methods, a robot positions a closed container close to one or more microphones and then shakes the container. This shaking motion induces collisions among the granules inside the container and between the granules and the container walls, producing a sound pattern that varies according to material properties such as grain size and softness. For instance, Eppe et al. [eppe2018deep] used a humanoid robot to shake closed capsules filled with different materials, collected the audio recordings, and utilized a neural network to classify the samples. Although the study did not focus exclusively on granular materials, the majority of the ten materials examined can be considered granular. Similarly, Jin et al. [jin2019open] used a robot arm to rotate capsules filled with granular materials by 90 or 180 degrees and recorded audio samples from the interactions. The audio samples were then pre-processed by a noise reduction algorithm and transformed into the Mel-Frequency Cepstral Coefficient space to perform classification.
@@
Vibration-based methods either complement data from microphones with other sensor modalities or use a completely different approach for vibration sensing. The work by Chen et al. [chen2016learning] is similar to audio-based methods since the work utilizes recordings from two microphones and a 3-axis accelerometer. The authors recorded data from 12 different materials that were sealed in containers that were shaken by a robot. However, the 12 materials differed significantly from each other, and only a few of them can be considered granular, with the rest being solid objects like an L-bracket or a piece of acrylic. Wang et al. [wang2023stev] showcased their novel electronic skin vibration sensor by classifying granular materials in a closed bottle. The authors used their soft robot gripper to shake the bottle and achieved high classification accuracy using different machine-learning models for the recorded signals, which were transformed from the time domain to the Fourier domain.
From a general viewpoint, we argue that the shaking-a-closed-container interaction is suboptimal when the classification information is used for manipulation-related downstream tasks. This is because the indirect interaction barely gives information about the material’s friction and flowing resistance properties, which are highly important for planning downstream manipulation actions such as scooping. Even though we are neither trying to estimate the friction parameters or flowing resistance explicitly, our proposed method directly interacts with the materials so that these properties are captured in the data. Thus, our method bases its classification decisions directly on the properties that affect the material’s manipulation-time dynamics the most.
@@ III Interactive Material Identification
The proposed interactive material identification framework consists of three parts: interactive exploration, feature extraction, and classification. At the interactive exploration stage, the robot interacts with the materials directly and collects feedback with an F/T sensor. The feature extraction stage is used to process the raw F/T samples to enhance their classifiability. Finally, we employ a classification model based on error-correcting output codes (ECOC) and support vector machines (SVM) to identify the materials.
III-A Interactive Exploration
The interactive exploration stage involves the robot interacting with materials directly in an open container. The robot interacts with the materials with a cylinder-shaped end effector. The six-axis F/T sensor is attached between the robot’s wrist and the end effector, as depicted in Figure 1(a). In this setup, we utilize a 7-DoF Franka Emika Panda robot arm along with an ATI Mini45 F/T sensor. The F/T sensor measures data at a sampling frequency of 500 Hz throughout the motion.
To maintain a standardized exploration procedure, the height of the material surface and the robot starting position are predetermined and maintained consistently. The robot initiates its interaction with the material from the designated starting position, ensuring that the end effector is nearly fully submerged in the material, as depicted in Figure 1(b). We employ a Cartesian velocity controller to guarantee constant velocity regardless of the materials. The exploration trajectory entails linear motion along the x-axis at a velocity of 0.03 m/s for 3.5 seconds. The recording of F/T data is started a fraction of a second before the robot executes the exploration trajectory to ensure that the information-rich beginning of the interaction is fully recorded.
Although the F/T sensor in the robot records data for six axes, we focus only on the forces along the x-axis for classification purposes (the open dataset contains full data). This is because the movement of the robot’s end effector is limited to the x-axis direction, making the information measured along this axis the most significant. This decision simplifies the classification task by reducing the input dimensionality for the classification model.
@@ III-B Feature extraction
To enhance the classifiability of the materials, we designed a feature space that takes into account their distinguishing characteristics, which are further discussed in Section IV. Specifically, we introduce a feature space representation called high-frequency magnitude histogram (HFMH) of the signal. In short, HFMH is the magnitude histogram of the F/T signal after high-pass filtering. We use a histogram with 100 bins evenly spaced in the range [-1.5, 1.5]. The range was defined to correspond to the maximum and minimum magnitudes present in the high-pass filtered F/T signals. For high-pass filtering, we use the Butterworth filter of order eight [proakisdigital], designed with respect to the desired cutoff frequency. Formally, we design a linear system defined by the transfer function
$H(z)=\frac{b_{0}+b_{1}z^{-1}+b_{2}z^{-2}+...+b_{8}z^{-8}}{1+a_{1}z^{-1}+a_{2}z^{-2}+...a_{8}z^{-8}}$ (1)
so that once the input signal is processed with the system, frequencies that are lower than the desired cutoff frequency get attenuated while higher frequencies are preserved. The best cutoff frequency, 23 Hz, was experimentally defined.
The designed feature space, neglecting the low-frequency content in the measurements, suggests that the materials are best separable based on the measurements’ high-frequency content, or lack thereof. The feature space also preserves information on the signal magnitudes to some extent, even though part of it is lost when the low frequencies are filtered out. The use of histograms leads to a compact representation with the cost of some information loss on the signal dynamics, that is, how the signal develops from the beginning to the end.
@@ III-C Classification
We use the trinary ECOC model [escalera2008decoding] for our multiclass classification task. The ECOC model divides the multiclass classification problem into a set of binary classification tasks and includes a coding and decoding scheme to determine the final class for the input. For a $K$-class classification task, we have $L=K\cdot(K-1)/2$ binary learners. Formally, the ECOC decision rule is then defined as:
[…]
@@
is the binary loss. Variable $s_{l}$ stands for the binary classifier score for the input sample.
We use SVM with linear kernel functions as binary classifiers. Linear kernels were experimentally found to be most suitable for our task. They also allow a fair comparison when we compare the proposed feature space representation to baseline feature spaces.
The decision to use ECOC and SVMs with manually designed feature extraction instead of a data-driven approach, such as a classification neural network, is driven by two main reasons. Firstly, the available data is sparse, making it difficult to train a data-hungry neural network. Additionally, the time-intensive nature of real-world data collection makes it impractical to simply increase the dataset size. Secondly, one of the main purposes of material identification is to support downstream manipulation tasks. A transparent identification method combined with a hand-crafted feature extraction increases the feasibility of integrating knowledge from material identification into planning and control algorithms compared to black-box data-driven approaches. Furthermore, our approach enhances the understanding of the problem, materials, and measured F/T signals themselves, laying the groundwork for further advancements in the field.
IV Dataset and data qualitative analysis
IV-A Dataset
We employed the exploration procedure described in section III to collect a dataset of 62 individual samples from each of the 11 materials. These materials span a range of everyday items, including food ingredients, gardening materials, and animal care supplies. The complete list of the materials is shown in Figure 2. The materials were chosen to reflect diversity in density, grain size, grain shape, and surface properties of the grains. To prevent material compression during the data collection, the materials were stirred in the container after each measurement iteration. Furthermore, for materials that are susceptible to grain breakage from repeated interaction with the end effector (such as potting gravel, clay granules, cat litter, and oat flakes), fresh material was introduced during the data collection.
We introduce two variations of the dataset in our experiments: the easy dataset and the hard dataset. The easy dataset contains only 8 materials and is designed so that the majority of the materials are supposed to be distinct from each other. However, even the easy dataset contains seemingly similar materials, such as wheat flour and finely ground breadcrumbs. The hard dataset consists of 11 materials, including more materials with similar characteristics. For example, the hard dataset includes two different kinds of clay granules (gardening clay granules and cat litter), along with fine sugar as a new addition to the fine-grained materials. The hard dataset is designed to enhance the challenge of the identification task, further evaluating the effectiveness of the proposed method.

### v2

@@ Abstract
Despite the potential the ability to identify granular materials creates for applications such as robotic cooking or earthmoving, granular material identification remains a challenging area, existing methods mostly relying on shaking the materials in closed containers. This work presents an interactive material identification framework that enables robots to identify a wide range of granular materials using only force-torque measurements. Unlike prior works, the proposed approach uses direct interaction with the materials. The approach is evaluated through experiments with a real-world dataset comprising 11 granular materials, which we also make publicly available. Results show that our method can identify a wide range of granular materials with near-perfect accuracy while relying solely on force measurements obtained from direct interaction. Further, our comprehensive data analysis and experiments show that a high-performancefeature space must combine features related to the force signal’s time-domain dynamics and frequency spectrum. We account for this by proposing a combination of the raw signal and its high-frequency magnitude histogram as the suggested feature space representation. We show that the proposed feature space outperforms baselines by a significant margin. The code and data set are available at: https://irobotics.aalto.fi/identify_granular/.
@@ I Introduction
The ability to accurately identify granular materials holds great potential across various domains of robotics, paving the way for future advancements in assistive robotics as well as autonomous working machines. In assistive robotics, the ability to distinguish different ingredients in food from each other is valuable to perform tasks such as assistive feeding or cooking assistance. Similarly, in industries such as mining or construction, autonomous work machines benefit from being able to differentiate between granular materials for tasks such as earthmoving or soil cultivation.
Most of the previous work on granular material identification or parameter estimation is done outside the robotic manipulation community, mostly in terramechanics and geoengineering, often utilizing tools such as Bevameter [apfelbeck2011systematic, kruger2023experimental] or cone penetrometer [white2022cpt, bian2025assessment]. Consequently, the approaches require specialized hardware with limited manipulability. This significantly limits the methods’ applicability in various domains of robotics and autonomous machines, where agility and multipurpose tools and sensors are often valued. In the robotic manipulation context, existing works mostly rely on tactile array sensors or audio sensing and identify materials based on their vibration signature when the material is shaken in a closed container [eppe2018deep, jin2019open, chen2016learning, wang2023stev]. However, in tasks such as earthmoving, the materials are not in closed containers, which limits the suitability of the methods based on vibration signatures.
To address this gap, we introduce a novel interactive perception framework that identifies granular materials by interacting with them directly. Specifically, our method measures feedback from the interaction with a standard 6-axis force-torque sensor (F/T sensor) mounted on the robot’s end effector, which is partially submerged in the material during a predefined motion. The collected F/T signals are subsequently classified based on the features extracted from the raw F/T measurements. We use custom-designed features based on our qualitative data analysis findings.
To the best of the authors’ knowledge, this work is the first attempt to identify granular materials using only a regular F/T sensor for perception. The contributions of this work include:
  • A novel interactive perception framework, enabling accurate identification of various granular materials, improving robots’ ability to understand and respond to the environment.
  • A novel real-world dataset of force-torque measurement sequences containing a total of 682 samples from 11 materials.
  • A comprehensive analysis of the dataset, facilitating a deeper understanding of its nature, the distinguishability of materials, and identification of their distinctive features.
@@ II-A Interactive perception with 6-axis F/T sensor
The majority of works on interactive perception with F/T sensors focus on identifying surface materials [liu12] or estimating surface properties [le2021probabilistic] of rigid bodies. Typically, these studies involve attaching the F/T sensor to an artificial fingertip or similar end-effector, which interacts with the material surface to explore its properties. For example,  [lam2014study] uses this approach to compare different classifiers, such as k-nearest neighbors method, support vector machines, and neural networks on their ability to identify 18 different surface materials. Similarly,  [markert2021fingertip] proposed an interaction procedure for an artificial fingertip, which first gathers F/T data from 21 different surface materials, then classifies the materials with high accuracy using features derived from the F/T measurements’ spectral distributions. Furthermore,  [le2021probabilistic] integrated F/T sensing with visual information to estimate friction maps for multi-material rigid objects.
In addition, F/T sensors have also been used for object recognition and property estimation for rigid objects. For example,  [higy2016combining] used 6-DoF F/T sensors combined with robot joint positions to perform haptic classification of rigid objects. In [pavlic2023robotscale], the authors proposed a method that employs a 6-axis F/T sensor attached to the robot’s two-fingered end effector to estimate the mass and inertia of rigid objects.
Prior research highlights that F/T sensors can provide valuable information for various robotic perception tasks. Our work expands the scope of F/T sensing by introducing a novel application area, granular materials, and specifically their identification.
@@ II-B Granular Material Identification
Alongside with liquid identification and property estimation [saal2010active, matl2019haptic, zhu2024liquids, elbrechter2015discriminating], research on identification of granular materials is sparse in the robotics literature. The problem is better studied in terramechanics and geoengineering, where a number of approaches have been proposed to characterize and identify different types of soil [apfelbeck2011systematic, kruger2023experimental, white2022cpt, bian2025assessment]. However, these methods rely on specialized and possibly large hardware, making them inapplicable for many robotic manipulation tasks. For example, the robot may have limited capacity to carry tools, making multipurpose tools and sensors a preferred option, or the task and the environment may require agile maneuvering that is not possible with physically large tools.
In robotics, the existing works on granular material identification have utilized audio recordings [eppe2018deep, jin2019open], vibration measurements [chen2016learning, wang2023stev], and capacitance sensing [hu2024robocaproboticclassificationprecision]. In addition to identification, parameter estimation of granular materials has been studied [matl2020inferring, guo2023estimating]. Furthermore, tactile sensor have been proposed for estimating the amount of granular materials and fluids in deformable containers [liu2023content].
In audio-based identification methods, a robot positions a closed container close to one or more microphones, shakes the container, and classifies the material based on the audio patterns created by colliding granules inside the container. For instance,  [eppe2018deep] used a humanoid robot to shake a closed capsule and used a neural network to classify the collected audio samples. Similarly,  [jin2019open] used a robot arm to rotate capsules filled with granular materials, recorded audio samples, pre-processed the samples by a noise reduction algorithm, and transformed them into the Mel-Frequency Cepstral Coefficient space to perform classification.
@@
Vibration-based methods complement microphone data with other sensor modalities or use specific vibration sensors. The method in [chen2016learning] is similar to audio-based methods since the work utilizes recordings from two microphones and a 3-axis accelerometer and classifies the recorded signals from a shaking action. In [wang2023stev], a novel electronic skin vibration sensor was introduced and applied successfully to recognize granular materials inside a closed bottle. A significant disadvantage of the vibration sensing-based methods is that they require either specialized hardware or a complicated sensor setup, which may be expensive and difficult to set up in dynamic real-world scenarios.
From a general viewpoint, we argue that the shaking-a-closed-container interaction is suboptimal when the classification information is used for manipulation-related downstream tasks. This is because indirect interaction barely gives information about the material’s friction and flowing resistance properties, which are highly important for planning downstream manipulation actions such as scooping or transportation. Further, in tasks such as earthmoving, the materials are not in closed containers, and thus the shaking approach is not directly applicable.
To summarize, our work aims to fill several gaps in the existing literature by proposing a granular material identification method that is not prone to background noise, can easily deal with materials that do not produce a loud sound when colliding, uses only a single, multipurpose sensing modality, thus requiring no specialized or complex hardware setups, and operates in the way that the interaction itself gives directly information about the dynamic properties of the material.
@@ III Interactive Material Identification
The proposed interactive material identification framework consists of three parts, as depicted in LABEL:fig:fig1: exploration, feature extraction, and classification. At the exploration stage, the robot interacts with the materials and collects measurements with an F/T sensor. The feature extraction stage is used to process the raw F/T measurements to enhance their classifiability. Finally, we employ a classification model based on error-correcting output codes (ECOC) and support vector machines (SVM) to identify the materials.
III-A Exploration
The exploration stage involves the robot interacting with materials directly in an open container. The robot interacts with the materials using a cylinder-shaped end effector, whose height is 80 mm and diameter 17 mm. The six-axis F/T sensor is attached between the robot’s wrist and the end-effector, as depicted in Fig. 1(a). In this work, we utilize a 7-DoF Franka Emika Panda robot arm and an ATI Mini45 F/T sensor. The F/T sensor measures data at a sampling frequency of $500\text{\,}\mathrm{Hz}$ throughout the motion.
To maintain a standardized exploration procedure, the height of the material surface and the robot starting position are predetermined and maintained consistently. The robot initiates its interaction with the material from the designated starting position, ensuring that the end effector is nearly fully submerged in the material, as depicted in Figure 1(b). We employ a Cartesian velocity controller to guarantee constant velocity regardless of the material. The exploration trajectory entails linear motion along the x-axis at a velocity of $0.03\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ for $3.2$ seconds. We decided to use a straight, one-way motion trajectory for simplicity, and our results show the simple approach is sufficient to capture the materials’ distinguishing characteristics. The choice on the velocity and duration of the motion was made based on our expectation on what is required to bring forth varying dynamic properties of different materials. The recording of the F/T data is started a fraction of a second before the robot begins the exploration trajectory to ensure that the information-rich beginning of the interaction is fully recorded.
@@ III-B Feature Extraction
To enhance the classifiability of the materials, we designed a feature space that considers their distinguishing characteristics, which are further discussed in Section IV. Specifically, we use the raw signal augmented by its High-Frequency Magnitude Histogram (HFMH), computed separately for each axis of the input F/T signal. In short, we define HFMH as the magnitude histogram of the F/T signal after high-pass filtering. We use a histogram with 100 bins evenly spaced in the range [1.5,1.5][-1.5,1.5]. The range was defined to correspond to the maximum and minimum magnitudes present in the high-pass filtered F/T signals. For high-pass filtering, we use the 8th orded Butterworth filter  [proakisdigital], designed with respect to the cutoff frequency of $23\text{\,}\mathrm{Hz}$. This cutoff frequency was found to be optimal by trial and error. All in all, the feature extraction stage results in 10200-dimensional feature vectors from which the raw time domain signal takes $9600$ and the HFMH representation $600$ dimensions.
@@ III-C Classification
We use the trinary ECOC model [escalera2008decoding] for our multiclass classification task. The ECOC model divides the multiclass classification problem into a set of binary classification tasks and includes a coding and decoding scheme to determine the final class for the input. For a $K$-class classification task, we have $L=K\cdot(K-1)/2$ binary learners. Formally, the ECOC decision rule is then defined as
$\hat{y}=\underset{k\in\{1,\ldots,K\}}{\mathrm{arg\,min}}\;\dfrac{\sum_{l=1}^{L}|m_{kl}|g(m_{kl},s_{l})}{\sum_{l=1}^{L}|m_{kl}|},$ (1)
@@
is the binary loss. Variable $s_{l}$ stands for the binary classifier score for the input sample. For a linear kernel SVM, $s_{l}$ is given by
$s_{l}=\mathbf{w}_{l}^{\top}\mathbf{x}+b_{l},$ (3)
where $\mathbf{x}$ is the input feature vector, and $\mathbf{w}_{l}$ is the weight vector given by the solution of the optimization problem
$\min_{\mathbf{w}_{l},b}\quad\frac{1}{2}\|\mathbf{w}_{l}\|^{2}$ (4)
$\text{s.t.}\quad y_{i}(\mathbf{w}_{l}^{\top}\mathbf{x}_{i}-b)\geq 1\quad\forall i\in\{1,\ldots,n\}$
where $y_{i}\in\{-1,1\}$ is the class label of $i$-th sample. Linear kernels were experimentally found to be most suitable for our task. They also allow a fair comparison when we compare the proposed feature space representation to baseline feature spaces.
The decision to use ECOC and SVMs with hand-designed feature extraction instead of an end-to-end data-driven approach, such as a classification neural network, was driven by two main reasons. Firstly, the time-intensive nature of real-world data collection makes it too costly to collect a sufficient dataset to train a data-hungry neural network. Secondly, transparent manually engineered features provide explainability and enhance the understanding of the problem, materials, and measured F/T signals themselves, laying the groundwork for further advancements in the field. Moreover, the explainability may facilitate the integration of the classification model into downstream planning and control algorithms.
IV Dataset and qualitative data analysis
IV-A Dataset collection
We employed the exploration procedure described in Section III to collect a dataset of 62 individual samples from each of the 11 materials, totaling 682 samples. We do not split these into fixed training and test sets, but instead we use repeated random partitions as we describe in V. These materials span a range of everyday items, including food ingredients, gardening materials, and animal care supplies. The entire set of the materials is shown in Fig. 2. The materials were chosen to reflect diversity in density, grain size, grain shape, and surface properties of the grains. However, also some similar materials, such as gardening clay granules and cat litter (which is also a type of clay granule), were selected to increase the difficulty of the identification task and to determine the bounds of the proposed method.
The materials were initially poured into a plastic container, while accurately measuring the desired amount by markers in the container. The material surface was smoothed before the first measurement and every subsequent iteration. To prevent material compression during the data collection, the material was carefully stirred in the container after each measurement iteration. Furthermore, fresh material was introduced during the data collection for materials susceptible to grain breakage from repeated interaction with the end effector (such as potting gravel, clay granules, cat litter, and oat flakes).